A garden measuring 6 meters by 8 meters is to have a pathway around it, increasing the total area to 252 square meters. What will be the width of the pathway?
Can you please explain HOW to find the answer? Please and thank you!

Answer by flame8855(424)

area of garden = 6*8 = 48 m^2
252-48= 204 m^2
length = 6 m
204/6= 34 m
so the width is 34 m